the equation of a line is y=7/8x+2
the slope fo a line paralled to it is ?
the slope of a line perpendicular to it is ?

Answer by mananth(16946)

compare with equation y= mx+b
where m is the slope

Parallel line slope m= 7/8
Perpendicular line slope = -8/7
( negative reciprocal of the slope of given line)